OpenStudy (anonymous):

x''+x' = 100sin5t Find a particular solution Xp(t) by using the complex function method

OpenStudy (fwizbang):

100 sin 5t = Im 100 e^{i5t}. So guess xp=Ae^{5it} plug it into the DEQ, solve for A, and then take Im xp.

OpenStudy (anonymous):

@fwizbang dude more inf please

OpenStudy (anonymous):

Can I reply :) .?!

OpenStudy (anonymous):

@Neemo of course :)

OpenStudy (anonymous):

Ok :) ! "A" here is a complex number ! \[xp'(t)=5iAe^{i5t}\] then \[xp"(t)=-25Ae^{i5t}\] so we should have \[(5Ai-25A)e^{i5t}=100e^{i5t}\]

OpenStudy (anonymous):

did you understand, ?should I continue ? @SkykhanFalcon

OpenStudy (anonymous):

@Neemo ok i understand

OpenStudy (anonymous):

then A= ???!

OpenStudy (anonymous):

@Neemo wait how about 5Ai ?

OpenStudy (anonymous):

sorry ! :) xp'(t)=5Aie^...did you undertand this ?

OpenStudy (anonymous):

@Neemo yeah i understand that i mean that (5Ai-25A) = 100 how can we know A

OpenStudy (anonymous):

\[5A(i-5)=100\] then \[A(i-5)=20\]

OpenStudy (anonymous):

oh well A is complex

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

dude something wrong i guess

OpenStudy (anonymous):

why ?!

OpenStudy (anonymous):

@neemo the answer should be (-10/13)*(5sin5t+cos5t)

OpenStudy (anonymous):

the book says so

OpenStudy (anonymous):

we haven't finished yet ! :) !

OpenStudy (anonymous):

so can you write A=x+iy ?

OpenStudy (anonymous):

of course i can @Neemo

OpenStudy (anonymous):

-(20i+100)/26

OpenStudy (anonymous):

It seems correct ! now can you compute Im(xp(t))=Im(Ae^i5t)

OpenStudy (anonymous):

@neemo how can i do that :/ i will be here in 20 minutes please :) something came up please wait :)

OpenStudy (anonymous):

kk I'll be here !don't worry !

OpenStudy (anonymous):

@Neemo ok here i am..you mean -20/26 = Ae^î5t ?

OpenStudy (anonymous):

back ! @SkykhanFalcon

OpenStudy (anonymous):

no ! \[xp(t)=Ae^{5 i t}=A(\cos(5t)+isin(5t))\] then compute Im(xp(t)) with A =-(20i+100)/26=-(10/13)i-(50/13)

OpenStudy (anonymous):

stiill there @SkykhanFalcon ?!

OpenStudy (anonymous):

ok i understand :) thank u so much :)

OpenStudy (anonymous):

Your're welcome anytime ! When you have an equation like x"+..x'+x=..cos(??t) then choose the Re(xp(t))

OpenStudy (anonymous):

ok i get it thank u again :)